OpenStudy (anonymous):

Write an equation of a line that is perpendicular to y = 2x + 3 and passes through (3,4).

OpenStudy (anonymous):

If the line is perpendicular, the slope of the original line times the slope of the new line has to be -1. So, the slope of the new line is -1/2.

OpenStudy (anonymous):

Then use the point-slope form to get \[y-4=\frac{ 1 }{ 2 } (x-3)\]

OpenStudy (anonymous):

i got positive 1/2

OpenStudy (anonymous):

The slopes have to multiply to -1/2

OpenStudy (anonymous):

One goes up and the other goes down :3

OpenStudy (anonymous):

how

OpenStudy (anonymous):

I mean -1 sorry not -1/2

OpenStudy (anonymous):

ok -1?

OpenStudy (anonymous):

So if the first slope is 2 and they have to multiply to -1 then the second is -1/2

OpenStudy (anonymous):

the second slope

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

Because \[-\frac{ 1 }{ 2 } * 2=-1\]

OpenStudy (anonymous):

Yes the second slope is -1/2

OpenStudy (anonymous):

So the slope of the perpendicular line is -1/2 because the slope of the first line is 2

OpenStudy (anonymous):

whats the first slope?

OpenStudy (anonymous):

The first slope is 2, because the first line is y=2x+3

OpenStudy (anonymous):

what would the equation be for the -1/2 slope/

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

Use the point-slope form to get y-4=-1/2*(x-3)

OpenStudy (anonymous):

and get y=-1/2x+11/2

OpenStudy (anonymous):

thanks i got the answer

OpenStudy (anonymous):

wait i got y=-1/2x+4

OpenStudy (anonymous):

I think you forgot to distribute the - sign when you multiplied -1/2 and -3